A boat takes 3.0 h to travel 36 km down a river, then 5.0 h to return. How fast is the river flowing? Express your answer in kilometers per hour.
Answer #1

Let the speed of the boat be x km/h

and speed of the river be y km/h

when it travels down the river

speed = distance/ time

(x + y) = 36 / 3  

x + y = 12 ----------(1)

when it tavels up stream

(x - y) = 36 / 5

x - y = 7.2 --------(2)

adding both equations we have

2x = 19.2

x = 9.6 km/h

y = 12 - x

= 12-9.6 = 2.4 km/h

so the speed of the river is 2.4 km/h